# Characterizing graph-nonedge pairs with single interval Cayley configuration spaces in 3-dimensions

### v1

@@ 1.2.1. Example
Finally, some of the concepts in this paper appear to be related to similarly long papers [12, 13] on generically globally-linked pairs in graphs. However, note that our study of $d$-convexity and $d$-Caley-connectedness does not require any genericity conditions.
1.3. Tools for working with graph minors.
Most work with minors is only concerned with proving that a graph either has or does not have a given minor, and the specifics of how this minor is obtained are not important. Hence, minors are often treated as equivalence classes of graphs. However, our results concern minors that are obtained without contracting a given edge. Furthermore, there is a natural minor map, formally defined below, between a graph $G$ and one of its minors $M$, and our proof techniques involve manipulating this map to show either that $M$ can be obtained while ensuring certain vertices of $G$ are mapped to specific vertices of $M$ or that some other graph $M^{\prime}$ is also a minor of $G$. Therefore, we need to introduce notation and terminology for working with minors in this manner, which we do below.
Figure 7. (A) a graph $G$, a subset $X\subseteq E(G)$ of edges contained in the black boxes, and a subset $Y\subseteq E(G)$ of dashed edges. (B) a minor $[G]$ witnessed by contractions $X$ and deletions $Y$. For the following, refer to Section 2.1. $H$ is the induced subgraph of $G$ shown in red and $[H]$ is the subgraph of $[G]$ induced by $[V(H)]$ shown in red. $J$ is the induced subgraph of $[G]$ shown in purple and $J^{-1}$ is the subgraph of $G$ induced by $\bigcup_{v\in V(J)}v^{-1}$ shown in purple. Any edge nonedge of $G$ whose endpoints are both contained in the same set $[u_{i}]^{-1}$ is contracted in $[G]$. The nonedge $u_{1}u_{2}$ is preserved and doubled in $[G]$ since $[u_{1}]\neq[u_{2}]$ and $[u_{1}u_{2}]^{-1}$ contains an edge. The nonedge $u_{2}u_{3}$ is retained in $[G]$ since $[u_{2}]\neq[u_{3}]$ and $[u_{2}u_{3}]^{-1}$ does not contain an edge.
We will only ever work with connected minors of connected graphs. Hence, assume $G$ and $M$ are connected. Then, $M$ can be obtained via some sequence of edge deletions and contractions, and no vertex deletions. From here on, all minors discussed are assumed to be obtained from such a sequence. Consider any sequence $G_{1},\dots,G_{n}$ of graphs such that $G=G_{1}$, $M=G_{n}$, and $G_{i+1}$ is obtained from $G_{i}$ via a single edge deletion or contraction. Also, consider the map from $V(G_{i})\cup E(G_{i})$ to $V(G_{i+1})\cup E(G_{i+1})\cup\{\emptyset\}$ that sends a deleted edge to the empty-set, the endpoints of a contracted edge to the vertex added to $G_{i+1}$ by this contraction, and every other vertex or edge to itself. By composing these maps, we get a map $\pi$ from $V(G)\cup E(G)$ to $V(M)\cup E(M)\cup\{\emptyset\}$, which we refer to as a minor map from $G$ to $M$. An edge $uv\in E(G)$ is contracted under $\pi$ if $\pi(u)=\pi(v)$ and deleted under $\pi$ if $\pi(uv)=\emptyset$. Contractions $X\subseteq E(G)$ and deletions $Y\subseteq E(G)$ witness minor $M$ if there exists a minor map from from $G$ to $M$ under which each edge in $X$ is contracted and each edge in $Y$ is deleted. For example, see to Figure 7. Note that all minors of $G$ witnessed by contractions $X$ and deletions $Y$ are isomorphic.
Next, assume contractions $X$ and deletions $Y$ witness $M$. We define a graph $[G]$ as follows. Let two vertices $u,v\in V(G)$ be equivalent if each edge along some path in $G$ between $u$ and $v$ is contained in $X$. Also, let $[u]^{-1}$ denote the equivalence class containing $u$. The vertex set $V([G])$ is the set of all such equivalence classes. For clarity, if we are referring to $[u]^{-1}$ as a vertex, then we will write it as $[u]$. Additionally, we write the pair $[u][v]$ as $[uv]$ and define the set $[uv]^{-1}$ to be $[u]^{-1}\times[v]^{-1}$ if $[u]\neq[v]$, and $[u]^{-1}$ otherwise. In the latter case, we call $[uv]$ a vertex of $[G]$ and treat it as such. The edge set $E([G])$ contains $[uv]$ if and only if $[u]\neq[v]$ and $[uv]^{-1}$ contains some edge in $E(G)\setminus Y$. Sometimes it will be convenient to refer to a vertex $x$ or an edge or nonedge $f$ of $[G]$ without making use of the bracket notation. The brackets are implied in these instances, and so $x^{-1}$ and $f^{-1}$ denote the sets $[x]^{-1}$ and $[f]^{-1}$, respectively. Note that $[G]$ is a minor of $G$, as it is clearly isomorphic to $M$. We use subscripts to differentiate minors denoted with brackets, e.g., $[G]_{A}$ and $[G]_{B}$ are two minors of $G$.
2. Contributions and proof obstacles
@@
From here on, we shorten “$d$-flattenability forbidden minor” to “$d$-forbidden minor,” and we omit $d$ from this terminology when it is clear from context. Given a minor $[G]$ of a graph $G$ and vertices $u,v\in V(G)$, define $uv$ to be contracted in $[G]$ if $[u]=[v]$ and preserved in $[G]$ otherwise. In the latter case, we call $[G]$ $uv$-preserving.
Theorem 4 (Main: $3$-SIP characterization).
@@
A graph-nonedge pair $(G,f)$ has the $3$-SIP if and only if no atom of $G\cup f$ that contains $f$ has an $f$-preserving $3$-forbidden minor.
The proof of Theorem 4 is presented in Sections 3 (converse direction) and 4 (forward direction). This theorem appears similar to a finite forbidden minor characterization, but has a few important differences. To discuss these differences, we must first extend the notions of minors and minor closed properties to graph-nonedge-set pairs. Given a graph $G$, a minor $[G]$, and a subset $X$ of vertices, edges, and nonedges of $G$, define the set $[X]=\{[x]:x\in X\}$. Letting $F$ be a set of nonedges of $G$ and $F^{\prime}$ be the set of nonedges in $[F]$, $([G],F^{\prime})$ is the minor pair of $(G,F)$ obtained via the sequence of edge deletions and contractions between $G$ and $[G]$. A property of graph-nonedge-set pairs is minor pair closed if the family of graph-nonedge-set pairs that have this property is closed under edge deletions and contractions.
In Example 1.2.1, it was shown that the family of graph-nonedge pairs with the $2$-SIP is not closed under edge deletions, and hence the $2$-SIP is not minor pair closed. Using a similar example based on $K_{d+2}$, it is easy to see that the $d$-SIP is not minor pair closed for any $d\geq 2$. Therefore, the $d$-SIP on the left-hand side of this theorem is not minor pair closed. On the other hand, the property “no atom of $G\cup f$ has an $f$-preserving $3$-forbidden minor” on the right-hand side of the theorem statement is minor closed. However, when we only consider atoms that contain $f$, this property is no longer closed under edge deletions. The intricacies of these statements and their differences are examined in Section 2.1.
2.0.1. Example
Theorem 4 shows that a graph-nonedge pair $(G,f)$ has the $3$-SIP if it is the pair in Figure 1(a), Figure 1(c), the top-left or bottom-left of Figure 2 such that $f$ is any of the dashed nonedges, Figure 3 such that $f$ is any of the dashed nonedges, Figure 6 left or right, Figure 8(a) or 8(b) such that $f$ is $w_{1}w_{2}$, or Figure 10(b) such that $f$ is any of the dashed nonedges. The same theorem shows that $(G,f)$ has the $3$-SIP if $G\cup f$ is any of the graphs in Figures 9(c), 9(e), 9(h), or 9(j), where $f$ is the top-most green edge. The fact that the pairs in Figure 8(a) or 8(b) have the $3$-SIP demonstrates that Theorem 2 does not extend to $d=3$ even after replacing “$K_{d+2}$ minor” with “$d$-forbidden minor” on the right-hand side. Theorem 4 also shows that $(G,f)$ does not have the $3$-SIP if it is the pair in Figure 1(d) or 4(a) or if $G\cup f$ is any of the graphs in Figures 9(a), 9(b), 9(d), 9(f), 9(g), or 9(i), where $f$ is the green edge.
Corollary 1 ($(d\leq 3)$-SIP characterization).
For any integer $d\leq 3$, a graph-nonedge pair $(G,f)$ has the $d$-SIP if and only if no atom of $G\cup f$ that contains $f$ has an $f$-preserving $d$-forbidden minor.
@@ 2.0.1. Example
For $d=3$, the theorem follows from Theorem 4. For any $d\leq 2$, by Theorem 2, it suffices to show that some atom of $G\cup f$ that contains $f$ has a $d$-forbbiden minor if and only if some atom of $G\cup f$ that contains $f$ has an $f$-preserving $d$-forbbiden minor. The converse direction is immediate and the forward direction follows from Theorem 5.2 in [22]. ∎
In Section 5, we conjecture that Corollary 1 is true for all $d\geq 4$. Corollary 2, below, follows immediately from Corollary 1.
Corollary 2 ($(d\leq 3)$-SIP is preserved in minors pairs of atoms-nonedge pairs).
@@
For any integer $d\leq 3$ and any graph-nonedge pair $(G,f)$ such that $G\cup f$ is an atom, if $(G,f)$ has the $d$-SIP, then so do all of its minor pairs.
As this corollary implies, the obstacle that causes the family $A^{d}$ of graph-nonedge pairs with the $d$-SIP to not be closed under edge deletions vanishes if we consider the subset of $A^{d}$ containing minor pairs of pairs $(G,f)$ such that $G\cup f$ is an atom. As we will see in Lemmas 1 and 2, below, $A^{d}$ is closed under contractions and a restricted type of edge deletion for any $d\geq 0$.
Figure 8. Graphs with $K_{5}$ and $K_{2,2,2}$ minors that can only be reached by contracting the edges labeled $w_{1}w_{2}$. See Proposition 1.
@@
Figure 9. See the above discussion, Examples 2.0.1, 4.0.1, 4.0.2, and 4.0.3, and Sections 2.1 and 4. (A), (B), (D), (F), (G), and (I) show graphs $G\cup f$ where $f$ is the green edge and $(G,f)$ is a minimal graph-nonedge pair (see Definition 3). Dashed green line-segments between vertices are nonedges. The dashed black boxes in (B), (D), (G), and (I) contain expanded $f$-components of top-level expanded $f$-separating cms’s of $G\cup f$, which are the purple and orange subgraphs in (B) and (D), the purple subgraph in (G), and the connected subgraph containing one purple edge and one blue edge in (I). (C), (E), (H), and (J) show minors of the graphs to their left obtained by contracting one purple edge in (C) and (H) and one blue edge in (E) and (J). The purple subgraphs in (C) and (H) and the blue subgraphs in (E) and (J) are $f$-separating cms’s whose $f$-components are the subgraphs in the dashed black boxes. The purple $f$-separating cms’s are in the top-level while the blue $f$-separating cms’s are not.
@@
Figure 10. (A) a $3$-tree $T$. (B) a $3$-connected partial $3$-tree obtained from $T$ by deleting the dashed edges.
@@
Next, Theorems 5 and 6, below, are notable byproducts of Theorem 4. Other notable byproducts of Theorem 4 are general lemmas about the structure of graphs with and without cs’s, which are collected in Appendix E. A $3$-tree is either $K_{4}$ or a graph obtained from a $3$-tree $T$ by adding a single vertex connected to each vertex in a clique in $T$ of size $3$. Equivalently, a $3$-tree is an edge-maximal tree-width $3$ graph. A partial $3$-tree is any subgraph of a $3$-tree. Equivalently, a partial $3$-tree is any graph that avoids having $K_{5}$, $K_{2,2,2}$, $V_{8}$, and $C_{5}\times C_{2}$ as minors [1]; see Figures 3 and 5. See Figure 10 for examples. Considerable effort has gone determining if a given set $F$ of nonedges of a partial $3$-tree $G$ is such that $G\cup F$ is a partial $3$-tree [7]. Theorem 5, which is proved in Appendix D, provides a new tool for this task. See Example 2.0.2, below. We write a countable indexed set $\{X_{1},X_{2},\dots\}$ is written as $\{X_{i}\}$.
Theorem 5 (Byproduct: $3$-connected partial $3$-tree star theorem).
@@
Proof sketch. First, we show that the theorem is true if $G\cup\{wu_{i}\}$ has no $3$-forbidden minor. Then, if this is not the case, we show that $G\cup wu_{i}$ has a $3$-forbidden minor for some nonedge $wu_{i}$. Since $G\cup wu_{i}$ is a partial $3$-tree, and hence has no $3$-forbidden minor, for each $wu_{i}$, this completes the proof.
@@
The above forbidden minor characterization for partial $3$-trees along with Theorem 3 shows that if a graph-nonedge-set pair $(G,F)$ is such that $G\cup F$ is a partial $3$-tree, then $(G,f)$ is $3$-convex. Theorem 6 shows that if $G\cup F$ is a partial $3$-tree, then $(G,F)$ has an even stronger property, which we now define. $(G,F)$ has the $d$-covering map property if, for any squared edge-length map $\ell$, the image of each connected component of the CS $\mathcal{C}^{d}(G,\ell)$ under the Cayley map $\phi_{F}$ is the entire CCS $\Omega^{d}_{F}(G,\ell)$.
@@
Theorem 6 (Byproduct: partial $3$-tree nonedge-set pairs have the $3$-covering map property).
Any graph-nonedge-set pair $(G,F)$ such that $G\cup F$ is a partial $3$-tree has the $3$-covering map property.
Proof sketch. First, observe that we can assume wlog that $G\cup F$ is a $3$-tree. For any squared edge-length map $\ell$, we show that the preimage of the interior of the CCS $\Omega^{3}_{F}(G,\ell)$ under the Cayley map $\phi_{F}$ can be partitioned into connected sets each of whose image under $\phi_{F}$ is the interior of $\Omega^{3}_{F}(G,\ell)$. Since $\phi_{F}$ is a closed map, the image of the closure of each of these sets under $\phi_{F}$ is $\Omega^{3}_{F}(G,\ell)$. Therefore, since each connected component $X$ of the CS $\mathcal{C}^{3}(G,\ell)$ must contain one of these sets, the image of $X$ under $\phi_{F}$ is $\Omega^{3}_{F}(G,\ell)$.
Next, we sketch how to obtain the above desired partition. Consider the set $A$ of $K_{4}$ subgraphs of $G\cup f$ each of whose points are not all coplanar in some $3$-realization of $(G,\ell)$. Let $\sigma$ be a vector with a coordinate $\sigma_{H}\in\{-1,+1\}$ for each $H\in A$. Consider the subset $\mathcal{C}^{3}_{\sigma}(G,\ell)$ of $\mathcal{C}^{3}(G,\ell)$ of realizations $p$ such that the sign on $\sigma_{H}$ matches the sign on the determinant representing the orientation of the points assigned to $H$ under $p$ for each $H\in A$. $\mathcal{C}^{3}_{\sigma}(G,\ell)$ is one of the connected sets that partition the preimage of the interior of $\Omega^{3}_{F}(G,\ell)$ under $\phi_{F}$, and there is one such set for each case of $\sigma$.
Finally, if $A$ is empty, then we show that $\mathcal{C}^{3}(G,\ell)$ is connected, and so the theorem is immediate. Otherwise, we make the observation that the preimage of each point in the interior of $\Omega^{3}_{F}(G,\ell)$ under $\phi_{F}$ intersects each set $\mathcal{C}^{3}_{\sigma}(G,\ell)$. Thus, showing that $\mathcal{C}^{3}_{\sigma}(G,\ell)$ is connected completes the proof.
2.0.2. Example
@@
Consider the partial $3$-tree $P$ in Figure 10(b) and let $F$ be the set of dashed nonedges. Since $P\cup f$ is a partial $3$-tree for any $f\in F$, Theorem 5 states that $P\cup F$ is a partial $3$-tree. Note that $P\cup F$ is the $3$-tree in Figure 10(a). Consider any squared edge-length map $\ell$ that assigns each edge of $P$ a non-zero value and such that the CS $\mathcal{C}^{3}(P,\ell)$ is non-empty. Also, let $u$ be the left-most vertex in Figure 10(b). For any realization $p$ in this CS, observe that the realization obtained by reflecting $p(u)$ across the plane defined by its neighbors is not in the same connected component of this CS as $p$. Even so, Theorem 6 states that the image of each connected component of this CS is the CCS $\Omega^{3}_{F}(P,\ell)$.
Finally, we present several independently interesting properties of $d$-convexity, $d$-Cayley-connectedness, and the $d$-SIP. In particular, Lemmas 1-3, below, are used throughout the proof of Theorem 4. See Example 2.0.3, below, for example applications of these results. A minor is induced if it can be obtained using only contractions. Note that induced minors and topological minors are not equivalent, i.e., while every topological minor is an induced minor, the converse is not true in general. A property of graph-nonedge-set pairs is induced minor pair closed if the family of graph-nonedge-set pairs with this property is closed under contractions.

### v2

@@ 1.2   Related Works
Finally, some of the concepts in this paper appear to be related to similarly long papers [26, 27] on generically globally-linked pairs in graphs. However, our study of $d$-SIP, $d$-convexity, $d$-Cayley-connectedness does not require any genericity conditions, and nonedges with $d$-SIP occur in sparser graphs overall than nonedges that are globally linked, weakly globally linked or even linked.
2   Contributions and proof obstacles
Note 2.
@@ 2   Contributions and proof obstacles
We shorten “$d$-flattenability forbidden minor” to “$d$-forbidden minor,” and we further omit $d$ from this terminology when it is clear from context.
Most usage of minors in the literature involves proving that a graph either has or does not have a given minor, and the specifics of how this minor is obtained are not important. Hence, minors are often treated as equivalence classes of graphs, as defined in Section 1.2. Instead, want to keep track of vertices as edge deletions and contractions are performed, and so we introduce rooted minors, which have been studied extensively [19, 28, 29, 30, 31, 32, 33]. Refer to Figure 10. Each vertex of a rooted minor of $G$ corresponds to a connected subgraph of $G$, all of whose edges have been contracted to obtain the minor. These connected subgraphs partition $V(G)$, hence, given a pair of distinct vertices of the rooted minor, their corresponding two connected subgraphs of $G$ have disjoint edge sets. A pair of vertices is an edge in the rooted minor only if the graph - induced by the union of the corresponding two connected subgraphs of $G$ is connected. In this case, the edge in the rooted minor corresponds to the set of edges of $G$ that have one endpoint in each of the corresponding two connected subgraphs. Note that the “if” direction of the previous statement does not hold. I.e., any edges straddling the two connected subgraphs could be deleted to obtain a nonedge in the rooted minor. A more formal definition of a rooted minor is given in Section 3. Unless otherwise specified, all minors in this paper are rooted. A pair of distinct vertices $u,v\in V(G)$ is contracted in a rooted minor if a vertex in the minor corresponds to a connected subgraph of $G$ containing both $u$ and $v$, and preserved in the minor otherwise. In the latter case, the minor is $uv$-preserving. The main result of the paper is a proof of Conjecture 1 in Section 1.2), restated as follows.
Theorem 4 (Main: $(d\leq 3)$-SIP characterization).
@@
For any dimension $d\leq 3$, a graph-nonedge pair $(G,f)$ has the $d$-SIP if and only if each atom of $G\cup f$ that contains $f$ has no $f$-preserving $d$-forbidden minor.
Remark 1.
This theorem is equivalent to Conjecture 1 since the right hand side can be stated equivalently as: “$f$ is contracted in each forbidden minor of each atom of $G\cup f$ that contains $f$.” For atoms $G\cup f$, the right hand side is even simpler: “$f$ is contracted in each forbidden minor of $G\cup f$.” However, inductive aspects of the proof - in both directions, even for the special case of atoms - appear to require the full strength of the general theorem, as discussed below.
We note that for any dimension $d\leq 2$, this result essentially follows from Theorem 2 and its proof in [12]. Specifically, the proof of Theorem 5.2 in [12], which is used to prove Theorem 2, already incorporates the additional ingredient that if $G\cup f$ has an atom that contains $f$ and has a $d$-forbidden minor, then it has an atom (in fact the same atom) with an $f$-preserving $d$-forbidden minor. Therefore our main contribution is for $d=3$, and the proof is presented in Sections 4 (converse direction) and 5 (forward direction). In Section 6, we conjecture that Theorem 4 is true for all $d\geq 4$.
Consider a natural approach to proving the forward “only if” direction of Theorem 4. Assume that when $G\cup f$ is a forbidden minor we have a certificate against $d$-SIP of $(G,f)$, i.e. a length assignment for the edges of $G$ for which there are two disjoint intervals of lengths attained by $f$. We can then extend this minor certificate to a certificate against $d$-SIP for general $(G,f)$ provided $f$ is retained in an induced forbidden minor of $G\cup f$. By induced minor, we mean a minor obtained only using edge contractions, without any edge deletions; and by retained in a minor, we mean that the pair $uv=f$ is not only preserved in a minor of $G$, but remains a nonedge. The certificate against $d$-SIP for $(G,f)$ is obtained by setting all lengths of contracted edges to 0 and the length of each non-contracted edge to that of its corresponding edge in the minor certificate. The following observation formalizes this approach.
Observation 1.
For any dimension $d\leq 3$, a graph-nonedge pair $(G,f)$ has the $d$-SIP only if $G\cup f$ does not have an $f$-retaining induced $d$-forbidden minor.
@@
Let $M$ be an $f$-retaining induced $d$-forbidden minor, and let $f^{\prime}$ be the nonedge in $M$ corresponding to $f$. The fact that $(M,f^{\prime})$ does not have the $d$-SIP follows relatively straightforwardly for $d\leq 3$ (see Example 1 and the proof of Lemma 19, if $M$ is $K_{d+2}$; and the proof of Theorem 8 in [17]; if $M$ is $K_{2,2,2}$). Along with Lemma 1, stated later in this section, it follows that $(G,f)$ does not have the $d$-SIP. ∎
To illuminate the obstacles to straightforward extensions of the forward “only if” direction of Observation 1 towards a proof of the forward direction of Theorem 4, we first inspect the relationships between a nonedge $f$ being preserved (resp. retained), in a minor (resp. induced minor). Observe that a nonedge $f$ is preserved in a minor $M$ of $G\cup f$ if and only if it is retained in $M$. One direction of this statement is immediate. For the other direction first delete the edges in $G$ whose endpoints correspond to $f$’s endpoints in $M$. While being preserved in a minor is equivalent to being preserved in an induced minor, the equivalence between preserved and retained is false for induced minors (where the minor is obtained using only contractions).
Due to this reason, Observation 1 cannot be extended beyond induced forbidden minors. In particular, Lemma 3 allows us to choose $(G,f)$ s.t. $(G,f)$ has $d$-SIP although $f$ is retained in a forbidden minor, i.e., preserved in an (induced) forbidden minor. This is achieved by some atom of $G\cup f$ that does not contain $f$ having a $d$-forbidden minor. For example, take $d=2$ and $(G,f)$ to be the pair on the left in Figure 6. In other words, this weaker property of $f$ being retained in a forbidden minor is insufficient as a certificate against $d$-SIP. However, if $G$ is an atom, the weaker property of $f$ being retained in a forbidden minor is a sufficient certificate for $(G,f)$ to not have $d$-SIP, for $d\leq 3$, as we show below in Corollary 1, of the forward direction of the main theorem applied to atoms.
From a larger perspective, Theorem 4 has the flavor of a finite forbidden minor characterization, although $d$-SIP is not minor closed, in particular, it is not closed under edge deletions. Specifically, only those atoms of $G\cup f$ that contain $f$ must avoid a certain presentation of forbidden minors; and atoms containing $f$ are allowed to have these forbidden minors as long as $f$ is contracted in them. For a more formal exploration, we introduce minor pairs of a graph-nonedge pair, analogous to minors of graphs. $(M,f^{\prime})$ is a (resp. induced) minor pair of a graph-nonedge pair $(G,f)$ if $M$ is an $f$-retaining (resp. induced) minor of $G$ and $f^{\prime}$ is the nonedge of $M$ corresponding to $f$. A property of graph-nonedge pairs is (resp. induced) minor pair closed if whenever a pair $(G,f)$ has the property, so do all its (resp. induced) minor pairs.
In Example 2, it was shown that the $2$-SIP is not closed under edge deletions, and hence this property is not minor pair closed. Using a similar argument, we see that the $d$-SIP (the property on the left-hand side of the theorem statement) is not minor pair closed for any $d\geq 2$. (Aside: later in this section, a repeatedly used Lemma 2 will show that the $d$-SIP is closed under a restricted type of edge deletion). However, Lemma 1 establishes that the $d$-SIP is induced minor pair closed. Not surprisingly, the right hand side of Theorem 4, namely the property “each atom of $G\cup f$ that contains $f$ has no $f$-preserving $d$-forbidden minor,” is closed under contractions but not under edge deletions, and therefore is induced minor pair closed, but not minor pair closed. However, if we strengthen this property to “$G\cup f$ does not have an $f$-preserving forbidden minor (*)” then it is closed under both edge deletions and contractions and therefore is minor pair closed. Moreover, when $G\cup f$ is an atom, this strengthened property (*) is precisely the right hand side of Theorem 4, and is necessary for $d$-SIP for $d\leq 3$. Hence, we obtain the following corollary for atoms.
We note that we are unaware of a direct proof of this corollary except as a consequence of the forward direction of Theorem 4: even for the special case of atoms, the induction in the forward direction appears to involve non-atoms.
Corollary 1 ($(d\leq 3)$-SIP is minor pair closed for atoms).
@@
For any dimension $d\leq 3$, the $d$-SIP is minor pair closed for atoms: for any graph $G$ with nonedge $f$, if $G\cup f$ is an atom and $(G,f)$ has the $d$-SIP, then all its minor pairs also have the $d$-SIP.
Next observe that the converse of Observation 1 does not hold even when $G\cup f$ is an atom. I.e. there are atoms $G\cup f$ where $(G,f)$ does not have $3$-SIP although $f$ is not retained in any induced forbidden minor. For example, let $G\cup f$ to be the graph in Figure 7(b) with $f$ being the solid green edge. But $f$ not being retained in any forbidden minor is precisely the strengthened property (*). And the converse of Theorem 4 shows sufficiency for $(G,f)$ to have $d$-SIP for $d\leq 3$ even for non-atoms (a core idea underlying the converse was illustrated in Example 1, Figures 1(a) 1(b), 1(c), and 12(d)). However, as in the case of Corollary 1 - i.e. the necessity of the property (*) for ($\leq 3$)-SIP for atoms - we are unaware of a direct proof of the sufficiency of (*) for ($\leq 3$)-SIP for atoms, except as a consequence of the general converse direction of Theorem 4: even for the special case of atoms, the induction in the converse direction appears to involve non-atoms.
Furthermore, as pointed out earlier, when $G\cup f$ is not an atom, property (*) is not necessary for $(G,f)$ to have $(\leq 3)$-SIP. In particular, Lemma 3 allows us to choose $(G,f)$ so that it has the $d$-SIP although some atom of $G\cup f$ that does not contain $f$ has a $d$-forbidden minor, e.g. the left Figure 6 for $d=2$,.
Next, Corollary 2, below, adds an apparently much weaker, but in fact, equivalent statement to Theorem 3 in Section 1.2 for dimensions $d\leq 3$.
Corollary 2 (Adding equivalent statement to Theorem 3 for $d\leq 3$).
For any dimension $d\leq 3$ and graph $G$ the following are equivalent.
  1. 1.
    $G$ is $d$-flattenable
  2. 2.
    for any subset $F\subseteq E(G)$, the pair $(G\setminus F,F)$ is $d$-convex
  3. 3.
    for each edge $f$ of $G$, the pair $(G\setminus f,f)$ has the $d$-SIP
Proof.
(1) $\implies$ (2) is Theorem 3. (2) $\implies$ (3) is immediate from definition of convexity. For (3) $\implies$ (1), assume to the contrary that (3) is true while (1) is false. Then, $G$ has a $d$-forbidden minor. Since each $d$-forbidden minor is an atom, some atom of $G$ must have a $d$-forbidden minor. Combining this with Corollary 3, below, allows us to assume wlog that $G$ is an atom. However, (3) and Theorem 4 imply that no edge of $G$ is preserved in any $d$-forbidden minor of $G$, which is a contradiction. ∎
@@
Figure 7: See Example 3 in this section, the discussion in Section 2.1, Examples 6-8 and the discussion in Section 5. Subfigures (a), (b), (d), (f), (g), and (i) show graphs $G\cup f$ where $f$ is the green edge and $(G,f)$ is a minimal graph-nonedge pair (Definition 3 in Section 5). Dashed green line-segments are nonedges. Top-level expanded $f$-separating CMSs of $G\cup f$ (Definition 4 in Section 5) are shown as the purple and orange subgraphs in (B) and (D), the purple subgraph in (G), and the connected subgraph containing one purple edge and one blue edge in (I). The expanded $f$-components of these CMSs are the subgraphs in the dashed black boxes. Contracting one purple edge in (b) and (d) yield (c) and (e), respectively. Contracting one bluw edge in (g) and (i) yield (h) and (j), respectively. The purple subgraphs in (c) and (h) and the blue subgraphs in (e) and (j) are $f$-separating CMSs whose $f$-components are the subgraphs in the dashed black boxes. The purple $f$-separating CMSs are in the top-level while the blue $f$-separating CMSs are not.
Example 3.
Theorem 4 can be used to verify the following statements. A graph-nonedge pair $(G,f)$ has the $3$-SIP if it is any of the pairs in Figures 1(a)-1(c), the top-left or bottom-left of Figure 2 such that $f$ is any of the dashed nonedges, Figure 3 such that $f$ is any of the dashed nonedges, Figure 6, Figure 8(a) or 8(b) such that $f$ is $w_{1}w_{2}$, and Figure 9(b) such that $f$ is any of the dashed nonedges. $(G,f)$ has the $3$-SIP if $G\cup f$ is any of the graphs in Figures 7(c), 7(e), 7(h), or 7(j), where $f$ is the top-most green edge. $(G,f)$ does not have the $3$-SIP if it is the pair in Figure 1(d) or 4(a) or if $G\cup f$ is any of the graphs in Figures 7(a), 7(b), 7(d), 7(f), 7(g), or 7(i), where $f$ is the green edge.
The fact that the pairs in Figure 8 have the $3$-SIP demonstrates that Theorem 2 does not extend to $d=3$ even after replacing “$K_{d+2}$ minor” with “$d$-forbidden minor” on the right-hand side. ∎
Figure 8: Graphs with $K_{5}$ and $K_{2,2,2}$ minors reachable only by contracting the edges labeled $w_{1}w_{2}$. See Example 3 in this section and the discussions in Sections 2.1 and 4. These figures are also referenced throughout the paper.
@@
Figure 9: (a) a $3$-tree $T$ and (b) a $3$-connected partial $3$-tree obtained from $T$ by deleting the dashed edges. See Example 4 and the discussion in this section.
@@
Next, Theorems 5 and 6, below, are independently interesting results concerning properties of partial 3-trees and covering maps. They are tools used in, and byproducts of, proving Theorem 4. They are proved in Section 4.2. A $3$-tree is either $K_{4}$ or a graph obtained from a $3$-tree $T$ by adding a single vertex connected to each vertex in a clique of size $3$ in $T$. Equivalently, a $3$-tree is an edge-maximal tree-width $3$ graph. A partial $3$-tree is any subgraph of a $3$-tree. Equivalently, a partial $3$-tree is any graph that avoids having $K_{5}$, $K_{2,2,2}$, $V_{8}$, and $C_{5}\times C_{2}$ as minors [34]; see Figures 3 and 5. See Figure 9 for examples. Considerable effort has gone into determining if a given set $F$ of nonedges of a partial $3$-tree $G$ is such that $G\cup F$ is a partial $3$-tree [35]. Theorem 5 provides a new tool for this task. See Example 4, below. For any integer $k\geq 1$, a graph is $k$-connected if it is connected and any of its separators has size at least $k$. We write a countable indexed set $\{X_{1},X_{2},\dots\}$ as $\{X_{i}\}$.
Theorem 5 ($3$-connected partial $3$-tree star theorem).
@@
The above-mentioned forbidden minor characterization for partial $3$-trees along with Theorem 3 shows that if a pair $(G,F)$ is such that $G\cup F$ is a partial $3$-tree, then $(G,F)$ is $3$-convex. Theorem 6 shows that if $G\cup F$ is a partial $3$-tree, then $(G,F)$ has an even stronger property, which we now define. $(G,F)$ has the $d$-covering map property if, for any squared edge-length map $\ell$, the image of each connected component of the CS $\mathcal{C}^{d}(G,\ell)$ under the Cayley map $\phi_{F}$ is the entire CCS $\Omega^{d}_{F}(G,\ell)$. Since $G\cup F$ is a partial $3$-tree, the result implies that the CS $\mathcal{C}^{d}(G,\ell)$ is a so-called branched covering space of the base space $\Omega^{d}_{F}(G,\ell)$ given by the Cayley map $\phi_{F}$. The $d$-covering map property further excludes the existence of lower dimensional connected components of $\mathcal{C}^{d}(G,\ell)$ whose image under $\phi_{F}$ does not cover the base space.
@@
Theorem 6 (Partial $3$-tree nonedge set pairs have the $3$-covering map property).
For any graph $G$ and nonedge set $F$ where $G\cup F$ is a partial $3$-tree, $(G,F)$ has the $3$-covering map property.
Example 4.
@@
Consider the partial $3$-tree $P$ in Figure 9(b) and let $F$ be the set of dashed nonedges. Since $P\cup f$ is a partial $3$-tree for any $f\in F$, Theorem 5 states that $P\cup F$ is a partial $3$-tree. Note that $P\cup F$ is the $3$-tree in Figure 9(a). Consider any squared edge-length map $\ell$ that assigns each edge of $P$ a non-zero value so that the CS $\mathcal{C}^{3}(P,\ell)$ is non-empty. Also, let $u$ be the left-most vertex in Figure 9(b). For any realization $p$ in this CS, observe that the realization obtained by reflecting $p(u)$ across the plane defined by its neighbors is not in the same connected component of this CS as $p$. Even so, Theorem 6 states that the image of each connected component of this CS is the CCS $\Omega^{3}_{F}(P,\ell)$. ∎
Finally, we present several repeatedly used tool lemmas that capture $d$-convexity, $d$-Cayley-connectedness, and $d$-SIP. In particular, Lemmas 1-3 and Corollary 3, stated below, are used throughout the proof of Theorem 4. See Example 5, below, for example applications of these results. To state Lemma 1, we generalize the notion of a minor pair to sets of nonedges $F$ and $F^{\prime}$ of a graph $G$ and its minor $M$. We say $(M,F^{\prime})$ is a minor pair of $(G,F)$ if $M$ is a minor of $G$ and $F^{\prime}$ is a, possibly empty, set of nonedges of $M$ whose corresponding set of nonedges $F$ in $G$ are retained in $M$.